Saturday 9 October 1999 9.00am EST
|During the last 24 hours Chandra exited the radiation belts, completed a post-belt CTI measurement and completed the GTO observation of EO102-72.3 with ACIS-S and the HETG. The GTO observation of AB Dor with ACIS-S and HETG was started at 7AM 10/9. The spacecraft systems continued to operate nominally.
Analysis of the ACIS CTI data from the radiation passage on day 275 shows that no further degradation in the energy resolution of the front-side illuminated CCD chips has occurred.
Return to Status Reports